# LongVLM: Efficient Long Video Understanding via Large Language Models

### v1

@@ Abstract
Empowered by Large Language Models (LLMs), recent advancements in VideoLLMs have driven progress in various video understanding tasks. These models encode video representations through pooling or query aggregation over a vast number of visual tokens, making computational and memory costs affordable. Despite successfully providing an overall comprehension of video content, existing VideoLLMs still face challenges in achieving detailed understanding in videos due to overlooking local information in long-term videos. To tackle this challenge, we introduce LongVLM, a straightforward yet powerful VideoLLM for long video understanding, building upon the observation that long videos often consist of sequential key events, complex actions, and camera movements. Our approach proposes to decompose long videos into multiple short-term segments and encode local features for each local segment via a hierarchical token merging module. These features are concatenated in temporal order to maintain the storyline across sequential short-term segments. Additionally, we propose to integrate global semantics into each local feature to enhance context understanding. In this way, we encode video representations that incorporate both local and global information, enabling the LLM to generate comprehensive responses for long-term videos. Experimental results on the VideoChatGPT benchmark and zero-shot video question-answering datasets demonstrate the superior capabilities of our model over the previous state-of-the-art methods. Qualitative examples demonstrate that our model produces more precise responses for long videos understanding. Code is available at https://github.com/ziplab/LongVLM.
@@ 1 Introduction
Large language models (LLMs) [36, 37, 44, 45, 42, 7] have revolutionized natural language understanding tasks and have demonstrated a remarkable capability to follow human instructions and intentions, emerging as a universal agent for general-purpose assistants. Drawing from the development of LLMs, Multi-modal Large Language Models (MLLMs) [30, 64, 8, 15] have driven the advance in vision-language learning by integrating visual encoders with LLMs and finetuning on language-image instruction-following data. However, developing Video-based Large Language Models (VideoLLMs) still poses a significant challenge due to the necessity of processing a large number of tokens for jointly modeling spatial-temporal dependencies across consecutive video frames. For instance, employing OpenAI CLIP-ViT-L/14 [40] as a visual encoder for a 100-frame video clip necessitates handling 25.6K visual tokens, leading to impractical computational costs with existing LLMs. To address this, recent approaches propose to extract video representation via precompression over visual tokens, utilizing pooling operation [35, 34] or query aggregation [24, 60] over the video token sequence before feeding them into the LLM, as shown in Fig. 1(a). While these models showcase impressive capabilities in providing a meaningful understanding of video content, they still face challenges in achieving a significant advantage in fine-grained understanding of long-term videos. For example, as shown in Fig. 1(b), while all models recognize the overall environment (workshop), the object (bike), and the action (fixing), previous methods may fail to correctly identify details such as the color of the clothes (brown), or the specific component being fixed (bicycle chain).
@@
The main reason is that long-term videos typically involve numerous key actions, complex activities, and camera movements. Consequently, a long video can be divided into a sequence of short-term segments. For instance, in the example depicted in Fig. 1(b), various short-term actions occur, e.g. speaking, displaying spare components, grabbing the bicycle chain, along with the camera moving from the human to the bicycle wheel, and eventually focusing on the broken chain. Similarly, prior methods in video recognition task suggest to decompose complex activities into sequences of sub-activities [46, 19, 14]. These approaches treat the features of each short-range activity as the local information within the videos and emphasize the importance of reasoning over local features to develop a temporal-structural understanding [46, 14, 47, 48, 19, 59] within long-term videos for comprehending fine-grained information. From this perspective, existing VideoLLMs treat all visual tokens equally and aggregate them into compact representations through pooling operations [35, 34] and query aggregation [24, 60]. While they successfully capture the global semantic context spanning the entire long-term videos, they often overlook preserving the local information for the short-term segments and the temporal structure of different short-term components, e.g., the order of events or sub-actions. However, exclusively modeling the temporal structure through the sequence of local features may still lead to inconsistent recognition across different segments and impede the overall understanding of the videos. To comprehend the content in long videos, the human visual system relies on a blend of local and global information [43]. Building on this insight, earlier approaches in video object detection [51, 53] suggest integrating global semantics into local localization descriptors, motivating us to include global semantic information into the sequence of local features for enriching the context understanding for each short-term segment.
@@ 2.2 Video-based Large Language Models
Building upon pretrained image-language models, traditional approaches [33, 13, 11, 22, 49, 28, 6, 26] are proposed to address various video-language learning tasks, e.g., video retrieval, video captioning, and video question answering. With the advent of LLMs, recent approaches are exploring the potential of developing Video-based Large Language Models (VideoLLMs) to unify various video-language understanding scenarios through human-video dialogue interactions. Existing VideoLLMs typically follow a common paradigm, which involves using a pretrained visual encoder to encode visual features, a projection layer to convert visual representations into the text latent space of LLMs, and a pretrained LLM for response generation. VideoChatGPT [35] and Valley [34] rely on pooling over visual tokens to obtain compact visual representations. VideoChat [24] utilizes pretrained video foundation models [25, 49] and Q-Former from BLIP-2 [23] to aggregate video representations. Video-LLaMA [60] proposes a Video Q-Former and an Audio Q-Former, enabling multiple modalities for video comprehension, while Video-ChatCaptioner [5] employs ChatGPT [36] to summarize video descriptions in multiple rounds of interactive question-and-answer conversation. Multiple video-centric instruction datasets [34, 24, 35] have also been proposed to finetune VideoLLMs for better video understanding capacity. Moreover, BT-Adapter [31] proposes a temporal adapter alongside the visual encoder for post-pretraining, while Video-Teller [29] highlights the importance of modality alignment in pretraining. Overall, these VideoLLMs rely on pooling and query aggregation on the whole long videos to extract visual representation for developing VideoLLMs, which lack of local information for fine-grained understanding in long videos. In contrast to existing VideoLLMs, we propose a simple yet effective framework that is feasible for aggregating both local and global information in long-term videos and preserves fine-grained content understanding.
@@ 3.2 Local Feature Aggregation
After obtaining the frame-level patch feature $\{{\bf P}^{t}\}_{t=1}^{T}$ from the last second layer of visual encoder, previous methods either apply factorized spatio-temporal pooling [35, 34], or utilize query aggregation [24, 60] over all visual tokens, which may miss local information referring to the short-term events or actions. Nevertheless, videos have heavy spatio-temporal redundancy, which results in redundant computational costs by directly considering all the patch features as the local representation for each segment. Therefore, we propose to aggregate compact visual features within each short-term segment. Specifically, we collect the patch feature for the $s^{th}$ segment ${\bf V}^{s}=\{{\bf P}^{t}\}_{t=(s-1)K}^{sK}\in{\mathbb{R}}^{KN\times d}$ and apply a hierarchical token merging module to aggregate the local feature while reducing the number of visual tokens. Inspired by ToMe [1], we resort to the bipartite soft matching method and gradually merge the visual tokens for each short-term segment. At the $i^{th}$ step, we randomly partition the $R_{i}$ tokens into two non-overlap token sets ${\mathbb{P}}_{i}$ with $r_{i}$ tokens and ${\mathbb{Q}}_{i}$ with $R_{i}-r_{i}$ tokens, where initial $R_{0}=KN$. Then we calculate the similarity scores between the tokens in set ${\mathbb{P}}_{i}$ and ${\mathbb{Q}}_{i}$ based on the patch features. To obtain the similarity scores, each visual token is divided into $C$ heads along channel dimension, each with $\frac{d}{C}$ channels. The similarity score for each token pair is obtained by averaging the cosine similarity scores over all heads following

### v2

@@ Abstract
Empowered by Large Language Models (LLMs), recent advancements in VideoLLMs have driven progress in various video understanding tasks. These models encode video representations through pooling or query aggregation over a vast number of visual tokens, making computational and memory costs affordable. Despite successfully providing an overall comprehension of video content, existing VideoLLMs still face challenges in achieving detailed understanding in videos due to overlooking local information in long-term videos. To tackle this challenge, we introduce LongVLM, a straightforward yet powerful VideoLLM for long video understanding, building upon the observation that long videos often consist of sequential key events, complex actions, and camera movements. Our approach proposes to decompose long videos into multiple short-term segments and encode local features for each local segment via a hierarchical token merging module. These features are concatenated in temporal order to maintain the storyline across sequential short-term segments. Additionally, we propose to integrate global semantics into each local feature to enhance context understanding. In this way, we encode video representations that incorporate both local and global information, enabling the LLM to generate comprehensive responses for long-term videos. Experimental results on the VideoChatGPT benchmark and zero-shot video question-answering datasets demonstrate the superior capabilities of our model over the previous state-of-the-art methods. Qualitative examples demonstrate that our model produces more precise responses for long videos understanding. Code will be available at https://github.com/ziplab/LongVLM.
@@ 1 Introduction
Large language models (LLMs) [36, 37, 45, 46, 43, 7] have revolutionized natural language understanding tasks and have demonstrated a remarkable capability to follow human instructions and intentions, emerging as a universal agent for general-purpose assistants. Drawing from the development of LLMs, Multi-modal Large Language Models (MLLMs) [30, 65, 8, 15] have driven the advance in vision-language learning by integrating visual encoders with LLMs and finetuning on language-image instruction-following data. However, developing Video-based Large Language Models (VideoLLMs) still poses a significant challenge due to the necessity of processing a large number of tokens for jointly modeling spatial-temporal dependencies across consecutive video frames. For instance, employing OpenAI CLIP-ViT-L/14 [40] as a visual encoder for a 100-frame video clip necessitates handling 25.6K visual tokens, leading to impractical computational costs with existing LLMs. To address this, recent approaches propose to extract video representation via precompression over visual tokens, utilizing pooling operation [35, 34] or query aggregation [24, 61, 42] over the video token sequence before feeding them into the LLM, as shown in Fig. 0(a). While these models showcase impressive capabilities in providing a meaningful understanding of video content, they still face challenges in achieving a significant advantage in fine-grained understanding of long-term videos. For example, as shown in Fig. 0(b), while all models recognize the overall environment (workshop), the object (bike), and the action (fixing), previous methods may fail to correctly identify details such as the color of the clothes (brown), or the specific component being fixed (bicycle chain).
@@
The main reason is that long-term videos typically involve numerous key actions, complex activities, and camera movements. Consequently, a long video can be divided into a sequence of short-term segments. For instance, in the example depicted in Fig. 0(b), various short-term actions occur, e.g. speaking, displaying spare components, grabbing the bicycle chain, along with the camera moving from the human to the bicycle wheel, and eventually focusing on the broken chain. Similarly, prior methods in video recognition task suggest to decompose complex activities into sequences of sub-activities [47, 19, 14]. These approaches treat the features of each short-range activity as the local information within the videos and emphasize the importance of reasoning over local features to develop a temporal-structural understanding [47, 14, 48, 49, 19, 60] within long-term videos for comprehending fine-grained information. From this perspective, existing VideoLLMs treat all visual tokens equally and aggregate them into compact representations through pooling operations [35, 34] and query aggregation [24, 61, 42]. While they successfully capture the global semantic context spanning the entire long-term videos, they often overlook preserving the local information for the short-term segments and the temporal structure of different short-term components, e.g., the order of events or sub-actions. However, exclusively modeling the temporal structure through the sequence of local features may still lead to inconsistent recognition across different segments and impede the overall understanding of the videos. To comprehend the content in long videos, the human visual system relies on a blend of local and global information [44]. Building on this insight, earlier approaches in video object detection [52, 54] suggest integrating global semantics into local localization descriptors, motivating us to include global semantic information into the sequence of local features for enriching the context understanding for each short-term segment.
@@ 2.2 Video-based Large Language Models
Building upon pretrained image-language models, traditional approaches [33, 13, 11, 22, 50, 28, 6, 26] are proposed to address various video-language learning tasks.With the advent of LLMs, recent approaches are exploring the potential of developing Video-based Large Language Models (VideoLLMs) to unify various video-language understanding scenarios through human-video dialogue interactions. Existing VideoLLMs typically follow a common paradigm, which involves using a pretrained visual encoder to encode visual features, a projection layer to convert visual representations into the text latent space of LLMs, and a pretrained LLM for response generation. VideoChatGPT [35] and Valley [34] rely on pooling over visual tokens to obtain compact visual representations. VideoChat [24] utilizes pretrained video foundation models [25, 50] and Q-Former from BLIP-2 [23] to aggregate video representations. Video-LLaMA [61] proposes a Video Q-Former and an Audio Q-Former, enabling multiple modalities for video comprehension, while Video-ChatCaptioner [5] employs ChatGPT [36] to summarize video descriptions in multiple rounds of interactive question-and-answer conversation. Recently, MovieChat [42] proposes an effective memory management mechanism to enable LLMs to reason over hour-long videos. Multiple video-centric instruction datasets [34, 24, 35, 42] have also been proposed to finetune VideoLLMs for better video understanding capacity. Moreover, BT-Adapter [31] proposes a temporal adapter alongside the visual encoder for post-pretraining, while Video-Teller [29] highlights the importance of modality alignment in pretraining. Overall, these VideoLLMs rely on pooling and query aggregation on the whole long videos to extract visual representation for developing VideoLLMs, which lack of local information for fine-grained understanding in long videos. In contrast to existing VideoLLMs, we propose a simple yet effective framework that is feasible for aggregating both local and global information in long-term videos and preserves fine-grained content understanding.
@@ 3.2 Local Feature Aggregation
After obtaining the frame-level patch feature $\{{\bf P}^{t}\}_{t=1}^{T}$ from the last second layer of visual encoder, previous methods either apply factorized spatio-temporal pooling [35, 34], or utilize query aggregation [24, 61, 42] over all visual tokens, which may miss local information referring to the short-term events or actions. Nevertheless, videos have heavy spatio-temporal redundancy, which results in redundant computational costs by directly considering all the patch features as the local representation for each segment. Therefore, we propose to aggregate compact visual features within each short-term segment. Specifically, we collect the patch feature for the $s^{th}$ segment ${\bf V}^{s}=\{{\bf P}^{t}\}_{t=(s-1)K}^{sK}\in{\mathbb{R}}^{KN\times d}$ and apply a hierarchical token merging module to aggregate the local feature while reducing the number of visual tokens. Inspired by ToMe [1], we resort to the bipartite soft matching method and gradually merge the visual tokens for each short-term segment. At the $i^{th}$ step, we randomly partition the $R_{i}$ tokens into two non-overlap token sets ${\mathbb{P}}_{i}$ with $r_{i}$ tokens and ${\mathbb{Q}}_{i}$ with $R_{i}-r_{i}$ tokens, where initial $R_{0}=KN$. Then we calculate the similarity scores between the tokens in set ${\mathbb{P}}_{i}$ and ${\mathbb{Q}}_{i}$ based on the patch features. To obtain the similarity scores, each visual token is divided into $C$ heads along channel dimension, each with $\frac{d}{C}$ channels. The similarity score for each token pair is obtained by averaging the cosine similarity scores over all heads following